OpenStudy (anonymous):

Complete. z2 + 8z – 20 = (z – 2)(z + __) -8 20 -10 10

OpenStudy (anonymous):

z^2 + 8z – 20 = (z – 2)(z + __) z^2 + 8z – 20 = (z – 2)(z +a) -2 + a = 8 -2*a = -20 so a =?

OpenStudy (triciaal):